OpenStudy (anonymous):

Find A^-1 in terms of A, where A^2-5A+2I=0.

OpenStudy (anonymous):

well, if \[A=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] then \[A^{-1}=\frac{2a}{-b\pm\sqrt{b^2-4ac}}\] ???

OpenStudy (anonymous):

These are matrices

OpenStudy (anonymous):

oh. K then \[ A(A-5)=-2I\\ \text{premultiplying with A^{-1}}\\ A^{-1}A(A-5)=-2A^{-1}I\\ A-5=-2A^{-1}\\ A^{-1}={1\over2}(A-5) \]

OpenStudy (anonymous):

@electrokid How do you subtract a matrix with a scalar?

OpenStudy (anonymous):

oh yea.. \[A^{-1}={1\over2}(A-5I)\]

OpenStudy (anonymous):

Does \(5A = A5I\)?

OpenStudy (anonymous):

interesting.

OpenStudy (anonymous):

I definitely need to brush up on my linear algebra.

OpenStudy (anonymous):

yes..